OpenStudy (anonymous):

A sample of college student GPA's are as follows: 2.5, 3.2, 3.8, 3.5, 2.8, 4.0 Find the mean absolute deviation. (Round your answer to the nearest hundreth.)

OpenStudy (kropot72):

\[\sum_{i=1}^{n}\frac{|x _{i}-\bar{x}|}{n}\]

OpenStudy (anonymous):

what it that

OpenStudy (anonymous):

step 1. find mean

OpenStudy (anonymous):

mean = (sum of all values) / number of values

OpenStudy (anonymous):

the average is 19.8

OpenStudy (anonymous):

so 19.8/6=3.3

OpenStudy (anonymous):

ok.. mean = 3.3

OpenStudy (anonymous):

correct so now what do I do with the 3.3

OpenStudy (anonymous):

the variance is 3.3 right

OpenStudy (kropot72):

Now you need to take each data value and subtract the mean. Then find the absolute value of the result. The first calculation is: 2.5 - 3.3 = -0.8 |0.8| = 0.8 Can you do the other 5 calculations and then add the 6 results? That result is then divided by 6 to find the Mean Absolute Deviation.

OpenStudy (anonymous):

ok so I got 0 as my next answer

OpenStudy (anonymous):

ooh.. wait. I was explaining the standard deviation.. follow @kropot72's response

OpenStudy (anonymous):

2. you subtract each number from the mean (i.e., from 3.3) and neglect the sign. then find the average of these values.

OpenStudy (anonymous):

I got 0

OpenStudy (kropot72):

Taking the second data value3.2: 3.2 - 3.3 = -0.1 |-0.1| = 0.1 Can you now calculate a result using the third data value 3.8?

OpenStudy (anonymous):

where does 3.8 come from

OpenStudy (anonymous):

idk what you mean

OpenStudy (kropot72):

Look at your question. 3.8 is the third data value from the left hand side.

OpenStudy (kropot72):

Can you subtract the mean (3.3) from 3.8: 3.8 - 3.3 = ?

OpenStudy (anonymous):

0.5

OpenStudy (anonymous):

ok so now what do I do

OpenStudy (kropot72):

The value you have just found is already positive. Therefore we do not need to find its absolute value. We just record 0.5 for future use. Taking the fourth value (3.5) just subtract 3.3 from it and record the result: 3.5 - 3.3 = ?

OpenStudy (kropot72):

@mbrousseau34 Are you there?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i got 00.6 as my answer

OpenStudy (kropot72):

Not really. The answer is found from \[\frac{0.8+0.1+0.5+0.2+0.5+0.7}{6}=?\]